Calculate the altman Z-score that is the question

12. Calculate the market-to-book ratio of the company as of Feb 2, 2019. This is the Feb 2, 2019 market capitalization of the2. Bankruptcy Risk. Calculate the Altman Z-Score as of the 2018 balance sheet date. To calculate the market value of equity,Consolidated Statements of Financial Position February 3, 2018 Ae Adivated $ 2.643 8,597 1,300 12,540 6.095 28,131 5,623 2,64Consolidated Statements of Operations 2017 $ 2018 74,433 $ 923 75,356 53,299 15,723 2,224 4,110 461 (27) 3,676 746 2,930 71.7Consolidated Statements of Comprehensive Income 2018 2,937 2017 As Adjusted (a) 2,914 2016 As Adjusted (a) 2,734 $ $ (million
Answer #1

Altman Z-score = (1.2 * working capital to total assets) + (1.4 * retained earnings to total assets) + (3.3 * EBIT to total assets) + (0.6 * market value of equity to total liabilities) + (1.0 * sales to total assets)

working capital = total current assets - total current liabilities = 12,540 - 13,052 = -512

total assets = 40,303

retained earnings = 6,495

EBIT = operating income = 4,110

market value of equity = price per share * shares outstanding = $68.97 * 517,761,600 = $35,710,017,552, which is $35,710 million

total liabilities = total current liabilities + total noncurrent liabilities = 13,052 + 15,600 = 28,652

sales = 74,433

Altman Z-score = (1.2 * working capital to total assets) + (1.4 * retained earnings to total assets) + (3.3 * EBIT to total assets) + (0.6 * market value of equity to total liabilities) + (1.0 * sales to total assets)

Altman Z-score = (1.2 * (-512 / 40,303)) + (1.4 * (6,495 / 40,303)) + (3.3 * (4,110 / 40,303)) + (0.6 * (35,710 / 28,652)) + (1.0 * (74,433 / 40,303))

Altman Z-score = 3.14

An Altman Z-score of 3 and above means that the company is unlikely to file for bankruptcy

No, the Altman Z-score does not suggest a high level of bankruptcy risk. It suggests a low level of bankruptcy risk.